caira12:

Solve for x: 7x + 15 = 38 (5 points) Select one: a. 23/7 b. 53/7 c. 23 d. 53

Vocaloid:

since you have 7x + 15, you can subtract 15 from both sides to isolate 7x what do you get after this step?

caira12:

?

Vocaloid:

7x + 15 - 15 = ?

caira12:

7x?

caira12:

oh wait

Vocaloid:

good, and since we subtracted 15 from the left side we must also subtract 15 from the right side 7x = 38 - 15 = 23

caira12:

yes 7x

Vocaloid:

so if 7x = 23 what does x = ?

caira12:

16?

Vocaloid:

good attempt but not quite since 7 is being multiplied to x, we would do the opposite of multiplying by 7, which is dividing by 7 so divide both sides by 7, what do you get?

caira12:

so divide 23 by 7 and 32 by 7?

Vocaloid:

hm, I'm not sure where you're getting 32 from you have 7x = 23 dividing both sides by 7 gives us 7x/7 = 23/7 x = 23/7 = your solution

caira12:

i meant 23 not 32 whoops haha